OpenStudy (anonymous):

The College of Business at UniTech is planning to begin an online MBA program. The initial start-up cost for computing equipment, facilities, course development, and staff recruitment and development is $350,000. The college plans to charge of $18,000 per student per year. However, the university administration will charge the college $12,000 per student for the first 100 students enrolled each year for administrative costs and its share of the tuition payments. a. How many students does the college need to enroll in the first year to break even? b. If the college can enroll 75 stud

OpenStudy (jack1):

stud...? wheres the rest dude?

OpenStudy (jack1):

@vishalini ?

OpenStudy (anonymous):

a. How many students does the college need to enroll in the first year to break even? b. If the college can enroll 75 students the first year, how much profit will it make? c. The college believes it can increase tuition to $24,000, but doing so would reduce enrollment to 35. Should the college consider doing this?

OpenStudy (jack1):

initial cost = -350000 admin charges = -12000x for x of up to 100 students college charges = +18000x x = number of students

OpenStudy (jack1):

y = profit y = gains - losses y = 18000x - 12000x(up to 100 max) - 350000 when y = 0, profits = losses ---> college breaks even

OpenStudy (jack1):

so when y = 0, x = ...? @vishalini ?

OpenStudy (jack1):

answer to (a) = solve below for x y = 18000x - 12000x(up to 100 max) - 350000 0 = 18000x - 12000x - 350000 answer to (b) = sub x = 75 into the below eqn, solve for y y = 18000x - 12000x - 350000 answer to (c) = sub x = 35 into below equation, solve for y, then compare answer with equation below that: y = 18000x - 12000x - 350000 where x = 35 VS y = 24000x - 12000x - 350000 where x = 35

OpenStudy (jack1):

that all make sense @vishalini ?

OpenStudy (anonymous):

thanks alot